Probably a Bad Question. It's been a while since high school,(like 20 years) but anyways. I have a 90 degree angle 10 foot in one direction 10 foot in the other. What is the formula used to find the distance between the 2 end points of the angle.
Answer by rfer(16322)

a^2+b^2=c^2
10^2+10^2=200
sqrt 200=14.14 ft